How to Get an Academic Position (or At Least, Try to Get One)

Lawrence Hubert, with a little help from his friends Where do you get the ad notices for what is open:

Two general sources: APS Observer (the best place; this is also online and keyword searchable); APA Monitor

Special Interest Groups (SIGs) that would vary from discipline to discipline. For Quantitative, for example, the Psychometric Society Web page, NCME, AERA Educational Researcher.

Also, things will get forwarded to you by your Division, for example What items are needed to apply for an academic position:

1) Research interests statement

2) Teaching interests statement

3) Resume

4) Cover letter (which usually can be brief if (1), (2), and (3) are good)

5) Reprints (or preprints of “in press” items)

6) Commitment from three extramural references to write for you — usually from your division and/or on your thesis committee.

For (5), provide a list of where you are applying, to whom the letter should be addressed, and what kind of position if is. They should have your items (1) through (4) as well before they write the letter.

Try not to have the same person write for different people applying for the same job.

(I have exemplars of items (1), (2), (3), and (4) from a few recent quantitative students)

You also need to prepare a “Job Talk” — usually, 50 minutes; and hopefully, one that will “knock their socks off”

When:

The application deadlines are in the Fall, and some maybe as early as October 1.

So, interviews may be scheduled before the middle of December; if not then, you have a break till, say, the second week of January.

Thus, you should have your proposal defense completed and passed by early October.

The academic ads appear mainly from August through November; after that, there may be some post-docs; industry or lab positions; temporary replacements; and the like.

The interview process, usually —

Arrive Sunday — picked up by, say, Chair of Search Committee

Monday/Tuesday — meet and greet faculty, students, and so on. Maybe see the facilities.

Leave Tuesday, late

On Monday, Job Talk for a general audience. Make it light and reasonably interesting, if you can; also, find a cute “hook”, if you can. Plan on 50 minutes, no more for the talk itself; questions ensue afterwards.

You may need to give one other talk, e.g., a specialized Quant talk (two theorems and a couple of lemmas); a Clinical talk in what would pass for the PSC here; teach a general class in social, say Usually, make your own travel arrangements and then get reimbursed. This may be so for the motel as well, but probably not for your meals. Never go on your own nickel!

And don’t do “if you take our offer, we will pay for your trip”, or “if we don’t make an offer, we will”. But if “we make an offer and you turn us down, it is on your nickel”

It is ridiculous how much emphasis is placed on the “Job Talk” — how come nobody remembers the unreliability of single behavioral instances

So, practice in your Brown Bags, and maybe for your friends as well.

When you meet and greet the usual faculty, they invariable ask if “you have any questions”

Some possibilities to fill the dead void:

What is it like to be female faculty member in the department (probably best asked of a women faculty member, and maybe only if you are a woman)

What is the usual “teaching load” — and to the Chair of the Search Committee, maybe — what might you expect me to take over.

What is the town like (“it is not the end of the world but you can see it clearly from here”)

What is the “climate” of the Department like. The import of this will vary with who you are.

For quant people, the question is: “do they just want a mule to do the teaching or are they really supportive of my research program”

How is the Department structured; how big. Number of students both undergraduate and graduate.

You will probably have an “exit” interview with the Department Head.

You both will “beat around the bush” with respect to salary and start-up. Now, good research oriented places start at about 65K.

If you have partner issues that will influence your decision, bring them up, but don’t “beat on them”. Explain what your partner does, and what he/she might be looking for.

The offer (if it comes) usually arrives in two parts —

A letter from the Dean that formally offers the job, usually with a salary figure only, and a reminder of not getting too comfortable until you get tenure (you cannot remain as an Assistant Professor beyond your seventh year — and the seventh year can be a terminal year).

Some “start-up” commitment letter from the Department Head.

See my example of what I sent to an incoming faculty member when I was Acting Head.

You will have some time to decide, but not a lot.

It’s not generally a good idea to pit offers against one another — even if you win, you go in with everyone hating you.

When you get a chance and can —

Go to job talks in the Department, even if not right in your area. See how it is done.

Prepare the required items, and update throughout your graduate career.

Publish, publish, publish — with your advisor, by yourself, and with other graduate students. You need to ”float to the top” of the search Committee’s applicant pool (but I won’t push this analogy further).

For Quant people, in particular, and for everyone else who wants to be respected, learn the basics of MATLAB in your area; also, learn the Latex/Tex system for your writing. Believe me, you will thank me later.

Some random observations —

Separate out your professional worth from your personal worth — this is the best advice I ever received (from Frank Baker when I began at Wisconsin). The process of review/publication and so on, is just brutal, and if you get everything tied up, you will be in trouble.

For a partner, try to find someone who is “portable”

— otherwise, see what spousal (or life partner) opportunities exist. Make sure you partner up with someone who is very good.

Dual career issues are always very difficult.

There are questions that cannot be asked during the interview — “are you married?”, “visa status”, ”are you going to get pregnant?”, etc.

See the handout, University of Illinois Guidelines for Preemployment Inquires